OpenStudy (anonymous):

HELP find the area of a circle whose circumference is 3.2 HELP a. 2.56 b. 10.24 c. 5.12

OpenStudy (anonymous):

circumference= 2*pi*r so, 2*pi*r=3.2 use this to find the radius and then use radius to find the area of circle

OpenStudy (anonymous):

You know that \[3.2=2 \pi r\] You can solve for r and plug into \[A=\pi r^2\]

OpenStudy (anonymous):

10.24?

OpenStudy (anonymous):

Not quite